# May the Feedback Be with You! Unlocking the Power of Feedback-Driven Deep Learning Framework Fuzzing via LLMs

### v3

@@ 5.2.1. Default fuzzing configuration
Figure 8. Code coverage comparison with SOTA baselines.
5.2.2. 24-hour fuzzing configuration
@@
Following prior studies (Xia et al., 2024; Yang et al., 2024), we run FUEL and baselines throughout a 24-hour fuzzing process. Table 2 reveals that FUEL generates the fewest tests (8.80% of TitanFuzz’s tests) on PyTorch. This phenomenon is attributed to our multi-agent architecture ($\mathcal{M_{A}}$ for feedback analysis and $\mathcal{M_{G}}$ for test generation) in the fuzzing loop, which is very time-consuming. Despite the huge computational overhead, FUEL shows competitive performance with TitanFuzz on code coverage. More importantly, FUEL demonstrates stronger bug detection ability as TitanFuzz fails to detect any new bugs on the latest version of PyTorch and TensorFlow. These results indicate that in DL framework testing, simple and extensive test generation (i.e., generating a single API) provides small rewards (especially in bug detection), emphasizing the need for strategic generation of architecturally diverse models rather than quantity-driven approaches. Regarding model validity, the near-perfect validity of TitanFuzz stems from its API-level fuzzing paradigm that inherently bypasses constraint verification between different operators. In the model-level testing approaches, FUEL demonstrates significant validity improvements over WhiteFox, achieving 24.82% and 41.63% higher validity rates in PyTorch and TensorFlow, respectively. This performance gap demonstrates FUEL’s enhanced capability in repairing invalid models through its self-repair strategy. The 24-hour experimental results position FUEL as the balanced solution that addresses the trade-off between model validity and diversity in DL framework fuzzing.
Table 2. Comparison with baselines over 24-hour fuzzing
@@
. Framework Technique # Bugs Coverage # Valid Tests (%) # Tests PyTorch FUEL 8 97,195 2,808 (78.28%) 3,587 TitanFuzz 0 98,487 36,734 (90.21%) 40,721 WhiteFox 1 88,390 3,944 (51.55%) 7,651 TensorFlow FUEL 2 94,114 1,930 (64.03%) 3,014 TitanFuzz 0 99,778 33,896 (93.84%) 36,121 WhiteFox 0 74,785 1,353(22.40%) 6,041
5.3. Bug Study
@@
Table 3 presents the bug detection results of FUEL over the last seven months. We first analyze some representative detected bugs and then analyze the advantage of FUEL to trigger similar new bugs. Lastly, we explain why some detected bugs won’t be fixed by DL framework developers.
Table 3. Summary of our detected bugs
. Frameworks Total Confirmed (Fixed) Pending Won’t Fix PyTorch 98 87 (49) 6 5 TensorFlow 6 6 (0) 0 0 Total 104 93 (49) 6 5
5.3.1. Analysis of bug examples
@@
As discussed in § 3.2, we use two oracles to identify bugs on two backends, including numerical inconsistency and behavior inconsistency. We introduce two representative bugs as follows.
@@
Figure 9. Representative bugs detected by FUEL
Figure 9(a) shows a numerical inconsistency bug of PyTorch which implements an incorrect graph-level optimization. The two tensors returned by torch.chunk are subtracted, resulting in different tensor outputs on the two backends. The root cause of this bug is that PyTorch Inductor attempts to fuse torch.nn.ReflectionPad3d, view, and torch.chunk in the computational graph (this optimization is also called operator fusion (Alwani et al., 2016)). However, in the final step of indexing (i.e., x[0]- x[1]), PyTorch Inductor incorrectly assumed that the value of the tensor must be positive (while it might be zero due to modulo operation), causing the compiler backend to output incorrect results. This bug was labeled as utmost priority by PyTorch community and assigned CVE-2025-46148, which was fixed by a PR immediately.
Figure 9(b) demonstrates a behavior inconsistency bug of TensorFlow, which is a security vulnerability detected by FUEL. The bug-inducing model consists of tf.reshape and tf.keras.layers.Conv2D. When setting padding='valid' for Conv2D, eager backend executes the model successfully, while compiler backend throws the negative dimension size error. This bug can be leveraged as a vulnerability by requiring the Conv2D parameter in padding='valid' format, resulting in DoS attacks, which was assigned CVE-2025-55559.
@@
Figure 10. The process of FUEL triggering new similar bugs with last triggered bugs
5.3.2. Analysis of detecting similar new bugs.
In our approach, one of the feedback goals is to detect new bugs with the last triggered bug. The insight behind this goal is that bug-inducing inputs are usually clustered together (Dickinson et al., 2001). $\mathcal{M_{A}}$ can analyze current bugs and guide $\mathcal{M_{G}}$ to generate similar bug-inducing inputs in the next iteration. FUEL can generate similar bug-inducing inputs from three perspectives, including operator overloading, parameter mutation, and inverse operators.
@@
“Thank you for the issues. Silent incorrectness is always high priority!” — A PyTorch core developer
@@
Inverse operators. Different from the above two perspectives, which trigger similar bugs on the same operator, FUEL can also target different operators. As Figure 10(c) shows, in the last execution, bitwise_right_shift outputs a different tensor on the compiler backend. $\mathcal{M_{A}}$ infers that an inverse operator (i.e., bitwise_left_shift) may have the same bug and suggests testing this operator in the next iteration. $\mathcal{M_{G}}$ subsequently generates the DL model consisting of bitwise_left_shift which triggers the silent incorrectness similar to bitwise_right_shift. This bug was labeled as high priority and assigned CVE-2025-46152.
5.3.3. Won’t Fix Bugs.
@@
5.4. Ablation Studies
Similar to the observations of prior studies (FuzzGPT (Deng et al., 2024b), WhiteFox (Yang et al., 2024)), we notice that the PyTorch community is more active than the TensorFlow community as PyTorch developers confirm and fix issues faster. Therefore, we run our ablation experiments on PyTorch with the default fuzzing configuration. Table 4 shows ablation comparison results.
Table 4. Comparison with ablation variants
@@ 5.4.2. Contribution of FASA
5.4.3. Contribution of different feedbacks
@@
6. Related Work
6.1. LLM-Based Fuzzing
@@
TitanFuzz (Deng et al., 2023) is one of the pioneering works in LLM-based fuzzing, which leverages two LLMs to generate and infill DL library APIs, respectively. Subsequently, Fuzz4All (Xia et al., 2024) believes LLMs have been trained on the multilingual corpus to serve as a universal fuzzer for the different SUTs with different programming languages. Sedar (Fu et al., 2024) leverages LLMs to generate high-quality initial seeds by adapting test cases from other Database Management Systems (DBMSs), improving code coverage in target DBMS fuzzing. LLMIF (Wang et al., 2024b) enhances IoT protocol fuzzing by using LLMs to extract protocol information and reason about device responses from protocol specifications. ChatAFL (Meng et al., 2024) utilizes LLMs to transform human-readable protocol specifications into machine-readable formats, enabling the fuzzer to generate and mutate protocol messages systematically. PromptFuzz (Lyu et al., 2024) employs coverage-guided prompt fuzzing techniques to generate fuzz drivers iteratively. More recently, KernelGPT (Yang et al., 2025) automates the generation of Syzkaller specifications using LLMs, leveraging their extensive pre-training on kernel code and documentation to create valid syscall sequences. Unlike previous LLM-based fuzzers, FUEL is the first fuzzing technique to leverage LLM to analyze various semantic information in the feedback and achieve promising performance.
6.2. DL Framework Testing
@@
In recent years, DL framework testing has become an active topic, and a large number of fuzzers have emerged, which can be classified into traditional fuzzers (Dutta et al., 2018; Pham et al., 2019; Wei et al., 2022a; Xie et al., 2022; Deng et al., 2022; Yang et al., 2023; Wang et al., 2020; Guo et al., 2020; Gu et al., 2022; Liu et al., 2023a, b; Mu et al., 2024) and LLM-based fuzzers (Deng et al., 2023, 2024b; Yang et al., 2024).
@@
One of the pioneering works is CRADLE (Pham et al., 2019), which detects inconsistencies by running pre-built DL models across multiple low-level backends of Keras (Keras, 2020). To enhance the diversity of models, LEMON (Wang et al., 2020) and AUDEE (Guo et al., 2020) build upon CRADLE by applying pre-defined mutation rules to seed models and inputs, generating a broader range of test cases. NNSmith (Liu et al., 2023a) models each DL operator with input constraints and shape inference, using SMT solving to generate valid models. Subsequently, NeuRI (Liu et al., 2023b) is proposed to synthesize DL models via rule inference automatically. Unfortunately, the above traditional fuzzers need manual design for mutation rules, operator specifications, or synthesis rules, which need domain-specific expert knowledge. With the success of LLMs in software testing (Lemieux et al., 2023; Feng and Chen, 2024), researchers have attempted to apply LLMs to fuzzing DL frameworks. TitanFuzz (Deng et al., 2023) is the first work to directly leverage LLMs to generate API call sequences. Then FuzzGPT (Deng et al., 2024b) is proposed to generate edge-case code via in-context learning and fine-tuning. WhiteFox (Yang et al., 2024) proposes a white-box compiler fuzzing technique aided by LLMs, targeting to generate specific inputs to trigger the optimization features in compilers. However, all of them struggle to generate diverse and valid DL models, and FUEL is proposed to fill this gap by adopting a heuristic search for operator selection and a program self-repair strategy by mining the potential of feedback.
7. Conclusion
@@
In this work, we propose FUEL, the first feedback-driven fuzzer for DL frameworks, which leverages LLMs to process feedback information in the fuzzing loop. FUEL employs an LLM-based multi-agent architecture comprising an analysis LLM and a generation LLM: the former infers concise analysis summaries from feedback information to guide the latter in generating tests that help to improve code coverage and bug detection. Moreover, based on multiple feedback guidance, FUEL adopts a heuristic search algorithm and program self-repair strategy to enhance model diversity and validity, respectively. Our extensive experiments demonstrate FUEL’s superiority over state-of-the-art DL framework testing techniques in both code coverage and bug detection.

### v4

@@ 5.3.1 Default fuzzing configuration
Figure 8: Code coverage comparison with SOTA baselines.
5.3.2 24-hour fuzzing configuration
@@
Following prior studies [xia2024fuzz4all, whitefox], we evaluate FUEL over a 24-hour fuzzing campaign. Table 2 reveals that FUEL generates the fewest tests (approximately 8.80% of TitanFuzz’s tests) on PyTorch. This phenomenon is attributed to the computationally intensive nature of the multi-agent architecture within our fuzzing loop. Despite the significant overhead, FUEL shows competitive code coverage compared to TitanFuzz and, more importantly, superior bug detection capabilities. Notably, TitanFuzz failed to detect any new bugs on the latest versions of PyTorch and TensorFlow. Similarly, NeuRI generates the most tests (588,610 on PyTorch and 649,924 on TensorFlow) by leveraging efficient inductive rule inference and its own engineering optimization. However, its reliance on random operator combinations tends to produce numerous simple or repetitive tests. Consequently, NeuRI detects fewer bugs and achieves lower code coverage than FUEL. These results indicate that in DL framework testing, simple and extensive test generation (i.e., generating a single API or random operator combination) provides small rewards, especially in bug detection, emphasizing the need for strategic generation of architecturally diverse models rather than quantity-driven approaches. In terms of model validity and diversity, the near-perfect validity of TitanFuzz stems from its API-level fuzzing paradigm that inherently bypasses the complex constraint verification required for operator interactions. Among model-level testing approaches, NNSmith achieves 100% validity rate by design as it utilizes SMT solvers to construct computation graphs. However, it relies on a small set of manually defined operators that are thoroughly tested, which restricts its ability to explore more unusual operators. FUEL demonstrates significant validity improvements over WhiteFox, achieving 24.82% and 41.63% higher validity rates in PyTorch and TensorFlow, respectively. This performance gap highlights FUEL’s enhanced capability in repairing invalid models through its self-repair strategy. The 24-hour experimental results position FUEL as the balanced solution that addresses the trade-off between model validity and diversity in DL framework fuzzing.
Table 2: Comparison with baselines over 24-hour fuzzing
@@
. SUT Technique # Bugs Coverage # Valid Tests (%) # Tests PT FUEL 8 97,195 2,808 (78.28%) 3,587 TitanFuzz 0 98,487 36,734 (90.21%) 40,721 WhiteFox 1 88,390 3,944 (51.55%) 7,651 NNSmith 1 87,922 103,248 (100.00%) 103,248 NeuRI 2 92,274 588,610 (93.09%) 632,310 TF FUEL 2 94,114 1,930 (64.03%) 3,014 TitanFuzz 0 99,778 33,896 (93.84%) 36,121 WhiteFox 0 74,785 1,353 (22.40%) 6,041 NNSmith 0 75,777 47,321 (100.00%) 47,321 NeuRI 0 76,493 649,924 (89.72%) 724,392
5.4 RQ2: Bug Study
@@
Table 3 presents the bug detection results of FUEL. We first analyze some representative detected bugs and then analyze the advantage of FUEL to trigger similar new bugs. Lastly, we explain why some detected bugs won’t be fixed by DL framework developers.
Table 3: Summary of our detected bugs
SUT Total Confirmed (Fixed) Pending Rejected
PT 98 87 (53) 6 5
TF 6 6 (0) 0 0
Total 104 93 (53) 6 5
5.4.1 Analysis of bug examples
@@
As discussed in § 3.2, we use two oracles to identify bugs on two backends, including numerical inconsistency and behavior inconsistency. We introduce two representative bugs from PyTorch and TensorFlow as follows.
@@
Figure 9: Representative bugs detected by FUEL
Figure 9(a) shows a numerical inconsistency bug of PyTorch which implements an incorrect graph-level optimization. The two tensors returned by torch.chunk are subtracted, resulting in different outputs on the two backends. The root cause of this bug is that PyTorch Inductor attempts to fuse torch.nn.ReflectionPad3d, view, and torch.chunk in the computational graph (this optimization is also called operator fusion [alwani2016fused]). However, in the final step of indexing (i.e., x[0]- x[1]), PyTorch Inductor incorrectly assumed that the value of the tensor must be positive (while it might be zero due to the modulo operation), causing the compiler backend to output incorrect results. This bug was labeled as utmost priority by PyTorch community and assigned CVE-2025-46148, which was fixed by a PR immediately.
Figure 9(b) demonstrates a behavior inconsistency bug of TensorFlow, which is a security vulnerability detected by FUEL. The buggy model consists of tf.reshape and tf.keras.layers.Conv2D. When setting padding='valid' for Conv2D, the eager backend executes the model successfully, while the compiler backend throws negative dimension size error. This vulnerability can be leveraged by requiring the Conv2D parameter in padding='valid' format, resulting in DoS attacks, which was assigned CVE-2025-55559. CVSS-3.x scored the severity of this vulnerability as "7.5 HIGH".
@@
Figure 10: The process of FUEL triggering new similar bugs with the last triggered bugs
5.4.2 Analysis of detecting similar new bugs
In our approach, one of the feedback goals is to detect new bugs with the last triggered bug. The insight behind this goal is that bug-inducing inputs are usually clustered together [dickinson2001pursuingfailure]. $\mathcal{M_{A}}$ can analyze current bugs and guide $\mathcal{M_{G}}$ to generate similar bug-inducing inputs in the next iteration. FUEL can generate similar bug-inducing inputs from three perspectives, including operator overloading, parameter mutation, and inverse operators.
@@ 5.4.2 Analysis of detecting similar new bugs
“Thank you for the issues. Silent incorrectness is always a high priority!” — A PyTorch core developer
@@
Inverse operators. Different from the above two perspectives, which trigger similar bugs on the same operator, FUEL can also target different operators. As Figure 10(c) shows, in the last execution, bitwise_right_shift outputs incorrectly on the compiler backend. $\mathcal{M_{A}}$ infers that an inverse operator (i.e., bitwise_left_shift) may have the same bug and suggests testing this operator in the next iteration. $\mathcal{M_{G}}$ subsequently generates the DL model consisting of bitwise_left_shift which triggers the silent incorrectness similar to bitwise_right_shift. This bug was labeled as high priority and assigned CVE-2025-46152.
5.4.3 Rejected Bugs
@@
5.5 RQ3: Ablation Studies
Similar to the observations of prior studies (FuzzGPT [fuzzgpt], WhiteFox [whitefox]), we notice that the PyTorch community is more active than the TensorFlow community as PyTorch developers confirm and fix issues faster. Therefore, we run our ablation experiments on PyTorch with the default fuzzing configuration. Table 4 shows ablation comparison results.
Table 4: Comparison with ablation variants
@@ 5.5.2 Contribution of FASA
5.5.3 Contribution of different feedbacks
@@
6 Discussion
6.1 Sensitivity Analysis of LLM Backends
To evaluate the sensitivity of FUEL on different LLM backends, we experiment with three combinations of $\mathcal{M_{A}}$ and $\mathcal{M_{G}}$. In addition to default implementation (§ 4), we consider two additional LLM pairs. (i) Qwen-2.5-Coder-32B [hui2024qwen2.5-coder] paired with Qwen3-30B [qwen3]: These two open-source LLMs were selected due to their compatibility with our local environment, and this setting allows us to examine how much performance degradation FUEL would experience in a zero-cost budget scenario. (ii) Claude-Sonnet-4 [claude-4] paired with GPT-5 [gpt-5]. These LLMs were chosen based on their SOTA performance on coding benchmark [swebench], and this setting allows us to investigate the performance gains when inspiring the maximum potential of FUEL.
[…]
Table 5: Performance of FUEL across different LLM backends
$\mathcal{M_{G}}$ $\mathcal{M_{A}}$ # Bugs Coverage \$Costs
Qwen-2.5-Coder-32B DeepSeek-V3 6 96,405 \$1.68
Qwen-2.5-Coder-32B Qwen3-30B 2 89,244 \$0.00
Claude-Sonnet-4 GPT-5 8 100,894 \$9.63
The above results indicate that FUEL is moderately sensitive to the choice of LLM backend. Lower-tier open-source LLMs tend to generate less diverse and valid tests and therefore detect fewer bugs. High-end frontier LLMs can better leverage feedback and thus improve the quality of both analysis and generation, although the improvements come at a substantially higher monetary cost. Such moderate sensitivity suggests that FUEL can maintain competitive performance even when deployed with mid-tier LLMs, while still retaining the ability to benefit from stronger LLMs when available. Notably, as future LLMs offer stronger reasoning, higher code generation, and more reliable feedback summarization ability, FUEL can directly inherit these advancements without requiring re-implementation. As a result, FUEL’s performance can be further enhanced with the progress of LLMs.
6.2 Limitations
FUEL incurs inherent computational costs due to the use of an LLM-based multi-agent architecture within the fuzzing loop, which results in notable challenges in both time and monetary overhead. (i) Time cost arises from the limited throughput of test generation in FUEL. As detailed in § 5.3.2, $\mathcal{M_{A}}$ must process a large amount of feedback in each iteration, which restricts the number of generated tests. Our current implementations demonstrate an average latency of 24.09 seconds per iteration. Fortunately, this limitation is expected to be alleviated by the rapid evolution of LLM inference optimization techniques [vllm]. (ii) monetary cost is introduced by the token-based billing model of commercial LLM API services, especially because feedback analysis consumes a substantial number of tokens. To reduce this expense, our implementation relies on cost-effective API providers such as the DeepSeek API [deepseek-api]. The complete experimental evaluation conducted using the DeepSeek API incurred a total cost of approximately \$40.89, corresponding to roughly \$0.39 per detected bug in our experiments.
7 Related Work
7.1 LLM-Based Fuzzing
@@
TitanFuzz [titanfuzz] is one of the pioneering works in LLM-based fuzzing, which leverages two LLMs to generate and infill DL library APIs, respectively. Subsequently, Fuzz4All [xia2024fuzz4all] believes LLMs have been trained on the multilingual corpus to serve as a universal fuzzer for the different SUTs with different programming languages. Sedar [fu2024sedar] leverages LLMs to generate high-quality initial seeds by adapting test cases from other Database Management Systems (DBMSs), improving code coverage in target DBMS fuzzing. LLMIF [wang2024llmif] enhances IoT protocol fuzzing by using LLMs to extract protocol information and reason about device responses from protocol specifications. ChatAFL [meng2024large] utilizes LLMs to transform human-readable protocol specifications into machine-readable formats, enabling the fuzzer to generate and mutate protocol messages systematically. PromptFuzz [promptfuzz] employs coverage-guided prompt fuzzing techniques to generate fuzz drivers iteratively. More recently, KernelGPT [kernelgpt] automates the generation of Syzkaller specifications using LLMs, leveraging their extensive pre-training on kernel code and documentation to create valid syscall sequences. Unlike previous LLM-based fuzzers, FUEL is the first fuzzing technique to leverage an analysis LLM to analyze various semantic information in the feedback and achieve promising performance.
7.2 DL Framework Testing
@@
In recent years, DL framework testing has become an active topic, and a large number of fuzzers have emerged, which can be classified into traditional fuzzers [pham2019cradle, wei2022free, xie2022docter, deng2022fuzzing, yang2023fuzzing, wang2020lemon, guo2020audee, gu2022muffin, liu2023nnsmith, liu2023neuri, mu2024devmut] and LLM-based fuzzers [titanfuzz, fuzzgpt, whitefox].
@@
One of the pioneering works is CRADLE [pham2019cradle], which detects inconsistencies by running pre-built DL models across multiple low-level backends of Keras [Keras2020]. To enhance the diversity of models, LEMON [wang2020lemon] and AUDEE [guo2020audee] build upon CRADLE by applying pre-defined mutation rules to seed models and inputs, generating a broader range of test cases. NNSmith [liu2023nnsmith] models each DL operator with input constraints and shape inference, using SMT solving to generate valid models. Subsequently, NeuRI [liu2023neuri] is proposed to synthesize DL models via rule inference automatically. Unfortunately, the above traditional fuzzers need manual design for mutation rules, operator specifications, or synthesis rules, which need domain-specific expert knowledge. With the success of LLMs in software testing [lemieux2023codamosa, feng2024prompting], researchers have attempted to apply LLMs to fuzzing DL frameworks. TitanFuzz [titanfuzz] is the first work to leverage LLMs to generate API call sequences directly. Then FuzzGPT [fuzzgpt] is proposed to generate edge-case code via in-context learning and fine-tuning. WhiteFox [whitefox] proposes a white-box compiler fuzzing technique aided by LLMs, targeting the generation of specific inputs to trigger optimization features in compilers. However, all of them struggle to generate diverse and valid DL models, and FUEL is proposed to fill this gap by adopting a heuristic search for operator selection and a program self-repair strategy by mining the potential of feedback.
8 Conclusion
@@
In this work, we propose FUEL, the first feedback-driven fuzzer for DL frameworks, which leverages LLMs to process feedback information in the fuzzing loop. FUEL employs an LLM-based multi-agent architecture comprising an analysis LLM and a generation LLM: the former infers concise analysis summaries from feedback information to guide the latter in generating tests that help to improve code coverage and bug detection. Moreover, based on multiple feedback guidance, FUEL adopts a heuristic search algorithm and program self-repair strategy to enhance model diversity and validity, respectively. Our extensive experiments validate the effectiveness of each component and demonstrate FUEL’s strong capability in detecting bugs and improving code coverage compared to SOTA baselines.
Acknowledgment
We thank Wei Cheng, Yanzhou Mu, and Zhiyuan Li for their insightful feedback on the manuscript.
Appendix A Failure Case Analysis
Figure 11: A failed self-repair of FUEL
As discussed in our ablation study, FUEL can repair most of the invalid models. However, not all the invalid models can be repaired. After our manual check, we find that failed repair attempts typically involve programs with excessive operators and complex constraints, where LLMs erroneously apply shape-conversion operators (e.g., View and Unsqueeze), zooming shape errors. Figure 11 presents a representative failure mode involving operator compatibility. The original model defines a Linar operator expecting an input dimension of 10, but receives a tensor x with shape (2, 50). Triggered by the dimension mismatch error, FUEL attempts to repair the invalid model by reshaping x into (10, 10) using additional .view(). While this satisfies the Linear operator’s constraint (producing an output shape (10, 20)), it implicitly destroys the batch dimension semantics. This “specious” fix causes a crash in the subsequent residual connection out + x, where the shapes (10, 20) and (2, 50) become irreconcilable. Finally, the patched model is still invalid, and both backends throw RuntimeErrors. A promising future direction is that LLM generates shape-agnostic DL programs (i.e., fuzz drivers) while leveraging constraint-solving techniques (e.g., Z3Prover [z3-solver]) to automatically resolve tensor shape and API parameters, thereby producing valid and diverse DL models through hybrid neuro-symbolic approaches.